OpenStudy (anonymous):

please help me with this *give an example of a rational function that is undefined at x=a where a is a real number. next state the domain of your rational function using both set and interval notation.

OpenStudy (anonymous):

can you help ?

OpenStudy (afrodiddle):

This is a rational function f(x) = 7/x-a

OpenStudy (anonymous):

thats all?

OpenStudy (afrodiddle):

I gave you an example. Make your own fill in the blanks :P \[f(x) = \frac{ blank }{ x-a }\]

OpenStudy (anonymous):

am i supposed to plug in numbers for a?

OpenStudy (afrodiddle):

Does it state so? then yes.

OpenStudy (afrodiddle):

the domain is all real numbers except a

OpenStudy (anonymous):

so i can have f(x)= 5/x-a?

OpenStudy (afrodiddle):

Yes.

OpenStudy (afrodiddle):

( - ∞ , a ) U ( a , + ∞ ) is the interval notation of a domain.

OpenStudy (anonymous):

and how can i get the interval and set notation?

OpenStudy (afrodiddle):

{ x ϵ ℝ | x ≠ a } Is the set notation.

OpenStudy (afrodiddle):

as I said, the domain is all real numbers except a.

OpenStudy (anonymous):

okay thanks so much!:)

OpenStudy (afrodiddle):

No problem.

OpenStudy (afrodiddle):

Dont forget to close the question